OpenStudy (anonymous):

sum or difference of Cubes: x^3 -8 sum or difference of Cubes: x^3 -8 @Mathematics

OpenStudy (anonymous):

this is a difference of cubes

OpenStudy (anonymous):

and it factors to... =(x-2)(x^2+2x+4)

OpenStudy (anonymous):

wait how does that work out....

OpenStudy (anonymous):

consider this: a^3+b^3=(a-b)(a^2+ab+b^2)

OpenStudy (anonymous):

sorry but that is confusing the hell out of me.... are you cubing the original.... is that allowed???

OpenStudy (anonymous):

remember that factoring is the opposite operation to multiplication; if you mutliply the right side of the equation that lagrange gave you (or the one that I gave you) you get the sum of cubes

OpenStudy (anonymous):

sums/difference factor according to a "formula" that you have to memorize; tere is no method like some other factoring

OpenStudy (anonymous):

i meant to write "sums/differences of cubes" on the last post

OpenStudy (anonymous):

i think i get it now...

OpenStudy (anonymous):

cool beans :})

OpenStudy (anonymous):

x^3-2^3 = (x-2)*(x^2 + 4x + 4)